OpenStudy (anonymous):

What is the product of the roots of x^2-4x+k =0 if one of the roots is 7?

OpenStudy (raden):

first, subtitute x = 7 into the equaiton : x^2-4x+k =0 (7)^2-4(7)+k =0 solve for k ?

OpenStudy (gorv):

product of roots=k sum of roots x1+x2=-(-4)/1=4 x1+x2=4 let x1=7 find x2 product=x1*x2

OpenStudy (anonymous):

K= -21?

OpenStudy (gorv):

yeah

OpenStudy (anonymous):

Then what do I do?

OpenStudy (gorv):

that is product of ur roots that's it